OpenStudy (anonymous):

Solve the system using elimination. 5x + 4y = 12 3x – 3y = 18

OpenStudy (anonymous):

I thought that's substitution i need a way using elimination....

OpenStudy (anonymous):

multiply the first equation by 3 and the second equation by 4. After you do that you should distribute. After distributing you can take out the 12y and -12y. And just solve from there.

OpenStudy (anonymous):

58 arisauotani Best Response 3After you figure out the x plug the x into one of the equation and solve for y

OpenStudy (anonymous):

5x + 4y = 12 3x – 3y = 18 First 3*(5x+4y)=3*12 4*(3x-3y)=4*18 then we get 15x+12y=36 12x-12y=72 15x+12y=36 +12x-12y=72 ___________________ 27x+0 =108 27x=108 x=4

OpenStudy (anonymous):

5x + 4y = 12 As x=4, so substitute the value of x into 5x + 4y = 12. 5*4+4y=12 20+4y=12 4y=12-20 4y=-8 Thus y=-8/4 y=-2 The solution of the equation is x=4 y=-2 Do I explain the question clearly? Do you understand now?

OpenStudy (anonymous):

Yes you do!

OpenStudy (anonymous):

I am glad to hear that.